an aqueous solution that contains 285mg of trichloroacetic acid (Cl3COOH) (M.wt=163g/mol), in 10ml (the acid is 73% ionized in water).
1
Expert's answer
2018-01-04T03:47:32-0500
Answer:
% = m solute / m solution
m (H 2 O) = ρ · V = 1 · 10 = 10 g
% (Cl 3 COOH) = m solute / m solution = (0.285 / (0.285 + 10)) · 100% = 2.771%